OpenStudy (anonymous):

Solve the following equation. Determine if the equation is an "identity" or a "contradiction." 2h + 4h - 5h = -3h + 4

OpenStudy (ghazi):

2h+4h-5h=-3h+4 6h-5h=-3h+4 h+3h=4 4h=4 h=4

OpenStudy (ghazi):

in the third line just add 3h on both the sides

OpenStudy (anonymous):

thanks

OpenStudy (ghazi):

:)